Closed Dome:
    • WFS powered up in the afternoon, cooled. no issues
8:30 Performing closed dome tests
  • Aligning to argos source. Aource appears vignetted on the ccd47.
    • FINAL Global offsets: 3.68, 3.33, -1.07, -86.86, 82.64, 0
  • We were off axis by 6mm on the baysides. once on axis we were no longer seeing the vignetting.
  • Closed through AOSGUI with WFS mag 9.8
  • Looks very turbulent. It was set to 1" seeing. Doug has changed the turbulence file in the daytime.conf on
    on wfs to point to ""
  • Apply optical gain with ratio 0.6,
  • DONEApply Optical gain is now available after "Done selected"
    ALERT!The Hold Gains still is missing the _v# in the path.
ESM closed dome tests - testing different frequencies
  • Optimize gain with ESM with brightest source 10.5 (WFS reading 9.9mag) bin1, frequency 300Hz gain: 6.48, 1.94 0.08 TN# 20201228_092938, average pupil intensity 160-180 photons/subapp

  • Setting Source to 10.6 on WFS
    • Clicked center star and it showed complete on FLAO gui but was offset
    • Illumination protection. We are getting bin 2, frequency 101, EM Gain 300 for mag 10.6 with TTM... strangeness.
    • Sending 14.6 mag, we get values we would expect: bin4 100Hz, mod 3 , emGain600. We are seeing 13.2mag, ~1500 average pupil intensity Adjusted the aperture on the argos source to 53 so that the estimated magnitude was closer to 14.6, ~500 photons/subapp.

    • Doug updated the table with an updated frequency of 15.5 to 300 Hz. The routine did an interposlation of the 14.6 mag target and determind the frequency to be 177Hz. We also noticed the bin4 reconstructor has 13 modes.
      Doug then updated the ACE table to have 300Hz for both 14.5 and 15.5. But the frequency is still reading 177Hz
    • Doug adjusted the frequncy to 400 for 14.5 and 15.5 mag bin 4 in teh ACE table and the returned frequency of our 14.6 star is 215Hz
    • Adjusted emGain from 600Hz to 400Hz but it did not seem to change when preset
  • Going to try sending ACE-AO preset to see if the interpolations in the ACE tables work in teh same way (or not the ways) we expect and try to see if we can figure out what is goin on with
  • Sending 16.6 mag bin4, we get the frequencies and gains that we expect
    • Sending preset for TTM 16.8mag. We should get 100Hz, emGain600 for both center pupils and close loop and we did.
      Sending preset for TTM 16.8 but his time we set Centering Loop frequency to 150 and the close loop freqency to 200. just showed us that we are getting what we would expect for bin 4. Sending again but this time interperolating between 16.5 (200Hz, 600gain) and 17.5 (100Hz, 600gain) and for 16.8 we get 125Hz...
    • Average pupil intensity 25-40, 7.6photons/subapp, estimated magnitude 17th mag at 125Hz.
    • Updated the flux column (the flux column impacted the frequency interpolations)
  • Preset for Estimated mag 17.1 TTMAO: Bin 4, frequency 200Hz, emGain 600 with no seeing disturbance: 0.9, 0.16, 0.1
    • ALERT!AUTO with TTM mode only inserts LN Gains and does not optimize or insert a reasonalbe gain file
  • Preset for Estimated mag 17.1 TTMAO: Bin 4, frequency 200Hz, emGain 600 with no seeing disturbance: 0.9, 0.16, 0.1, Average pupil intensity ~28 , 4.8 photons/subapp, TN# 20201228_120626. ( Disturbance file didnt take, so restarting arb, actually it had taken the gains were misleadingly low)
  • Restarting the WFS Arb, apparenly it did take.
  • Sending preset for 16.4mag (estimated mag 16.8 on WFS), bin 4, 211 Hz TN#20201228_122628, 20-30 average pupil intensity
  • Sending preset for 15.8mag (estimated mag 16.8 on WFS), bin 4, 319 Hz TN#20201228_124006, 45 average pupil intensity, 10.3 photons/subapp, safeskipping during optimize gain. Lots of vibrations of some sort.... Opening and resending preset. We are safeskipping TN#20201228_124818
  • AUXLoops died again
  • Resending preset to 15.8 but with a lower frequency, 200Hz. bin4, safeskipping continues. TN#20201228_125206
  • Resting and setting the shell, 640 modes
  • Resending TTM for 15.8 bin4, looks more stable although there is still quite a bit of tip tilt jitter TN#20201228_125833. ~50 Average pupil intensity, 10.1 photons per sub app, 15.7 estimated magnitude.

  • Sending Mag14.4 for bin2 400Hz test
    • Failed actuator alert 164
    • We didn't have any pupils during the close loop process.... the process claimed to be up, issue with pupil diamters. When I started taking TN#20201228_130924 the pupils magically reappeared. Camera lens railed in X right after. Resent preset
    • ~32 counts average pupil intensity, 6.7 photons per subapp. Camera lens continues to rail in X for bin 2 for 14.2 mag
  • Sending preset for 10.9mag bin2 400Hz TTM
    • 800-1000 average pupil intensity, 145 photons/subapp, estimated magnitude 10.9. camera lens railed again.
    • 10.5 mag has 215 photons/subapp, 1200-1400 average pupil intensity
  • Sending preset for 10.4mag bin1 412Hz. Estimated mag is 10.6. 55-70 counts average pupil intensity, 37.2 photons/subapp
    • Optimize gain was quick for bin4 TTM but takes forever for bin1
    • TN#20201228_1334444
  • Sending preset for 9.9 bin1 497 Hz
    • Center star claimed to be successful in finding star and moving it to the hotspot but star did not move: 13:38:08
    • Second try was successful, found star and moved it up to the hotspot TN#20201228_34315
    • 59.4 photons/subapp, ~100 counts average pupil intensity.
  • Could not park ARGOS cal unit in the morning. Would get to 56.68, 56.7, 56.71 but would not trigger the in park position. Asked David to remove stow pins and move down 5 deg in elevation but he could not enable elevation drive. David thought it was because the park position of the argos swing arm was not triggered. He replaced the stow pins and I tried to moving the swing arm to away from the park position and back in swveral times....It was very close but it would not trigger. David notified the day crew and said he would handle it. I shot Gustavo an email. I had to leave it

-- JenniferPower - 28 Dec 2020
Topic revision: r6 - 28 Dec 2020, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback